OpenStudy (anonymous):

\(x^2\) is a perfect square, right?

OpenStudy (tkhunny):

Typically. What brings it up?

OpenStudy (anonymous):

I am factoring a prefect square trinomial and I just wanted to make sure! :D So, thanks! @tkhunny

OpenStudy (anonymous):

Oh, quick question! is \(8x~=~or~\neq~\to~2(x^2 \times 4^2)\) @tkhunnу

OpenStudy (tkhunny):

?? \(2\cdot 4^{2} = 2\cdot 16 = 32\)